four macrostates of a system have 2,6,16 and 2 microstates respectively a physical quantity was measured be 1,3,4 and 7 in those state respectively find<x>
Answer 49/13
1
Expert's answer
2019-04-08T09:12:37-0400
"\left< x\right>=\frac{\sum x n_x}{\sum n_x}"

"\sum x n_x =(2)(1)+(6)(3)+(16)(4)+(2)(7)=98"

"\sum n_x =(2)+(6)+(16)+(2)=26"

"\left< x\right>=\frac{98}{26}=\frac{49}{13}"